# Rewarding Doubt: A Reinforcement Learning Approach to Calibrated Confidence Expression of Large Language Models

### v3

@@ 6 Results and Discussion
Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting. *Results are from the original paper [27].
@@
Verbalize 0.3459 0.5858 0.6310
Chain-of-Thought 0.3065 0.6379 0.6273
Top-K 0.1611 0.6673 0.6023
Surrogate Token 0.3686 0.5923 0.5933
Sequence Probability 0.3156 0.7804 0.5955
Self-Consistency 0.1134 0.8213 0.6224
LACIE* 0.1200 0.7200 n/a
Trained Probe 0.0189 0.8173 0.5925
Rewarding Doubt (ours) 0.0226 0.8592 0.6309
Calibration Performance
@@
To assess how well our approach improves calibration, we compare it against the zero-shot LLM baseline (Verbalize) and several established methods in both Single-Answer and Multiple-Answer question-answering tasks. Results for the Single-Answer setting on TriviaQA are presented in Table 1, and those for the Multiple-Answer setting on QAMPARI appear in Table 2. Across both tasks, Rewarding Doubt substantially improves the model’s confidence calibration over zero-shot verbalization.
@@
In the Single-Answer setting on TriviaQA, Rewarding Doubt achieves an ECE of 0.0226 and an AUROC of 0.8592, clearly outperforming all zero-shot baselines as well as LACIE, which is based on DPO-based optimization. The second fine-tuned method, Trained Probe, which relies on supervised fine-tuning, reports a slightly lower ECE (0.0189), both methods achieve near-perfect results. Further the AUROC of Rewarding Doubt is notably higher, suggesting that although both methods offer strong calibration, Rewarding Doubt better discriminates between correct and incorrect answers. In the Multiple-Answer setting on QAMPARI, Rewarding Doubt also outperforms baselines, achieving an ECE of 0.0816 and an AUROC of 0.6947. In comparison, Verbalize, Sequence Probability, and Trained Probe perform notably worse. Our findings support the claim by Azaria and Mitchell [3] that a model’s internal state encodes information about the truthfulness of statements, which can serve as an indicator of uncertainty. However, without fine-tuning, the model struggles to utilize this internal information effectively. Our approach enables the model to make use of this correlation and translate it into an accurate expression of the probability that a given answer is correct.
@@
The calibration curves in Figure 4 further illustrate these improvements. For both TriviaQA and QAMPARI, the fine-tuned model’s confidence much more closely aligns with the ideal 45° line than the zero-shot Verbalize baseline. Additionally, we observe a shift in the confidence distribution after fine-tuning. As shown in Figure 5, in a zero-shot setting the LLM (Verbalize) predominantly assigns high confidence scores (8 or above), reflecting overconfidence, a pattern also noted by Xiong et al. [37], who attribute it to supervised pretraining that favors confident expressions. After fine-tuning with Rewarding Doubt, the model’s confidence scores (shown in Figure 5(b)) span a wider range, including lower values, indicating a more nuanced expression of uncertainty. This shift suggests that fine-tuning mitigates overconfidence and better aligns the model’s confidence with its actual performance.
@@
Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting.
@@
Verbalize 0.5319 0.6047 0.2549
Sequence probability 0.5324 0.5942 0.1927
Trained probe 0.1117 0.6481 0.2233
Rewarding doubt (ours) 0.0816 0.6947 0.2480
@@
Table 3: Comparison of generalization results on the CommonsenseQA dataset trained on the TriviaQA dataset.
Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
Verbalize 0.2820 0.5425 0.6860
Trained Probe 0.4819 0.5374 0.7108
Rewarding doubt (ours) 0.2930 0.6385 0.7163
@@
To assess the generalization abilities of Rewarding Doubt, we evaluated the model trained on TriviaQA in out-of-domain settings using the CommonsenseQA [29] and MedQA [15] datasets. Results are shown in Table 3 and Table 4, respectively.
On MedQA, Rewarding Doubt significantly outperforms Verbalize in both metrics, while on CommonsenseQA, it achieves a comparable ECE, however paired with a much higher AUROC. This discrepancy highlights a limitation of relying solely on ECE for evaluating calibration. ECE does not reflect how well a model discriminates between correct and incorrect predictions across different confidence levels. A model consistently assigning moderate confidence values could appear well-calibrated under ECE, yet fail to offer meaningful distinctions between uncertain and certain cases. The AUROC, by contrast, directly measures this discriminative ability. Thus, the substantial improvements in AUROC underscore that Rewarding Doubt produces more useful and actionable confidence estimates. Compared to the Trained Probe baseline, Rewarding Doubt consistently outperforms with significantly improved metrics, showing a stronger ability to generalize to new datasets.
Table 4: Comparison of generalization results on the MedQA dataset trained on the TriviaQA dataset.
@@
Verbalize 0.4480 0.5075 0.5066
Trained Probe 0.2099 0.5513 0.5051
Rewarding doubt (ours) 0.1145 0.6649 0.5161
We also explore generalization across experimental settings by applying a model trained in a Single-Answer setting to a Multiple-Answer task. Although it does not reach the performance of a model trained specifically for that task, it still outperforms baseline methods considerably, demonstrating transferability of the learning confidence estimation patterns. This suggests promising applications for improving confidence estimation in more complex or less structured scenarios, such as fact verification and calibration in free-text generation, even when specialized training data is unavailable.
Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, while generalizing across tasks, and maintaining stable task performance, making it an effective approach for accurate confidence calibration and expression in language models. Beyond improvements in calibration quality, our method also offers practical advantages. While fine-tuning requires an initial training investment, inference remains highly efficient, as only a small, constant number of tokens need to be generated to express confidence. In contrast, zero-shot methods like Chain-of-Thought and Self-Consistency have substantial computational overhead during inference by requiring lengthy reasoning chains or multiple generations. Rewarding Doubt introduces no such overhead, does not rely on an additional model, and directly provides actionable confidence estimates through simple verbalization directly by the LLM, making it highly suitable for real-world deployment.
Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings.
@@
Base model Single fact 0.5875 0.5787
Single fact Single fact 0.1536 0.7240
Base model Multi fact 0.5319 0.6047
Single fact Multi fact 0.1777 0.6617
Multi fact Multi fact 0.1061 0.7268
@@
During our experiments, we observed that the policy occasionally converges to always predicting a fixed confidence value, likely caused by training instabilities. While this did not occur consistently, it suggests a potential area for further refinement, such as adjusting the reward structure or introducing additional constraints to encourage a more stable learning of a diverse confidence distribution. Additionally, our approach is currently limited to settings, where a binary correctness measure can be established. In the future this can be extended to also include continuous correctness measures such as usually applied for free text generation.
@@ Appendix A Prompts
For all the question-answering settings, the model is directly prompted to answer a question without a preceding example or context. For our method the model was prompted to answer the question and additionally provide a verbalized confidence. For the other baselines that do not need a verbalized confidence but infer it indirectly, the model is prompted to only give the correct answer. The specifics for multiple-choice are slightly changed but hold mostly the same meaning. The exact prompts for each method can be seen in Table 6 for open questions and Table 7 for multiple-choice questions. The prompts for each Multi-Answer method can be seen in Table 8. We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
@@ Appendix B Proof
$R(a,\hat{p},j)=\begin{cases}log(\hat{p}),&\text{if $j(a)$ = 1 (correct)}\\ log(1-\hat{p}),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
@@
is strictly negative for $\hat{p}\in(0,1)$. Hence, $f(\hat{p})$ is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
@@
$R(a,\hat{p},j)=\begin{cases}log(\max(\hat{p},\epsilon)),&\text{if $j(a)$ = 1 (correct)}\\ log(\min(1-\hat{p},1-\epsilon)),&\text{\text{if $j(a)$ = 0 (incorrect)}}\end{cases}$
@@
Appendix C Societal Impact
@@
Appendix D Asset Details
The TriviaQA dataset [16] is available at https://nlp.cs.washington.edu/triviaqa/ under the Apache License 2.0. The CommonsenseQA dataset [29] (https://www.tau-nlp.sites.tau.ac.il/commonsenseqa) and the MedQA dataset [15] (https://github.com/jind11/MedQA?tab=readme-ov-file) are licensed under the MIT License. The QAMPARI dataset [2] is published at https://samsam3232.github.io/qampari/ and licensed under the CC0 1.0 Universal license. All these licenses allow broad use, distribution and modification. We use the Meta-Llama-3-8B-Instruct [10] language model, provided at https://huggingface.co/unsloth/llama-3-8b-Instruct-bnb-4bit under the Llama 3 Community License Agreement.

### v4

@@ 5 Results and Discussion
Table 1: Comparison of methods on the TriviaQA dataset in the Single-Answer setting with 95% CIs in brackets. * Results are from the original paper (Stengel-Eskin et al., 2024) and include standard error.
@@
Verbalize 0.3459 [0.3375,0.3543] 0.5858 [0.5778,0.5936] 0.6310 [0.6222,0.6397]
Chain-of-Thought 0.3065 [0.2981,0.3157] 0.6379 [0.6284,0.6475] 0.6273 [0.6181,0.6363]
Top-K 0.1611 [0.1529,0.1695] 0.6673 [0.6580,0.6768] 0.6023 [0.5936,0.6110]
Surrogate Token 0.3686 [0.3595,0.3783] 0.5923 [0.5818,0.6027] 0.5933 [0.5844,0.6016]
Sequence Probability 0.3156 [0.3074,0.3237] 0.7804 [0.7725,0.7876] 0.5955 [0.5864,0.6040]
Self-Consistency 0.1134 [0.1066,0.1210] 0.8213 [0.8129,0.8298] 0.6224 [0.6131,0.6317]
PPO-M 0.3262 [0.3173,0.3346] 0.5274 [0.5227,0.5319] 0.5749 [0.5662,0.5835]
PPO-C 0.3607 [0.3524,0.3697] 0.5439 [0.5384,0.5491] 0.5258 [0.5164,0.5358]
LACIE* 0.1200 $\pm$0.02 0.7200 $\pm$0.02 n/a
Trained Probe 0.0189 [0.0147,0.0275] 0.8173 [0.8099,0.8250] 0.5925 [0.5834,0.6017]
Rewarding Doubt (ours) 0.0226 [0.0176,0.0302] 0.8592 [0.8523,0.8664] 0.6309 [0.6222,0.6399]
@@
To assess how well our approach improves calibration, we compare it against the zero-shot LLM baseline (Verbalize) and several established methods in both Single-Answer and Multiple-Answer question-answering tasks. Results for the Single-Answer setting on TriviaQA are presented in Table˜1, and those for the Multiple-Answer setting on QAMPARI appear in Table˜2. Across both tasks, Rewarding Doubt substantially improves the model’s confidence calibration over zero-shot verbalization.
@@
In the Single-Answer setting on TriviaQA, Rewarding Doubt achieves an ECE of 0.0226 and an AUROC of 0.8592, clearly outperforming all zero-shot baselines as well as LACIE, which is based on DPO-based optimization. The second fine-tuned method, Trained Probe, which relies on supervised fine-tuning, reports a slightly lower ECE (0.0189), both methods achieve near-perfect results. Further the AUROC of Rewarding Doubt is notably higher, suggesting that although both methods offer strong calibration, Rewarding Doubt better discriminates between correct and incorrect answers. In the Multiple-Answer setting on QAMPARI, Rewarding Doubt also outperforms baselines, achieving an ECE of 0.0816 and an AUROC of 0.6947. In comparison, Verbalize, Sequence Probability, and Trained Probe perform notably worse. Our findings support the claim by Azaria & Mitchell (2023) that a model’s internal state encodes information about the truthfulness of statements, which can serve as an indicator of uncertainty. However, without fine-tuning, the model struggles to utilize this internal information effectively. Our approach enables the model to make use of this correlation and translate it into an accurate expression of the probability that a given answer is correct.
@@
The calibration curves in Figure˜4 further illustrate these improvements. For both TriviaQA and QAMPARI, the fine-tuned model’s confidence much more closely aligns with the ideal 45° line than the zero-shot Verbalize baseline. Additionally, we observe a shift in the confidence distribution after fine-tuning. As shown in Figure˜5, in a zero-shot setting the LLM (Verbalize) predominantly assigns high confidence scores (8 or above), reflecting overconfidence, a pattern also noted by Xiong et al. (2024), who attribute it to supervised pretraining that favors confident expressions. After fine-tuning with Rewarding Doubt, the model’s confidence scores (shown in Figure˜5(b)) span a wider range, including lower values, indicating a more nuanced expression of uncertainty. This shift suggests that fine-tuning mitigates overconfidence and better aligns the model’s confidence with its actual performance.
@@
Table 2: Comparison of methods on the QAMPARI dataset in the Multiple-Answer setting with 95% CIs in brackets.
@@
Verbalize 0.5319 [0.5172,0.5461] 0.6047 [0.5837,0.6267] 0.2550 [0.2410,0.2698]
Sequence probability 0.5324 [0.5225,0.5432] 0.5942 [0.5775,0.6110] 0.1928 [0.1829,0.2024]
Trained probe 0.1117 [0.0997,0.1262] 0.6481 [0.6241,0.6726] 0.2233 [0.2094,0.2384]
Rewarding doubt (ours) 0.0816 [0.0723,0.0951] 0.6947 [0.6776,0.7113] 0.2480 [0.2348,0.2609]
@@
To test the consistency of our method across different models, we perform an ablation study across diverse LLM architectures and sizes. Specifically, we apply Rewarding Doubt to Qwen-2.5 (3B and 7B) and Gemma-2 (9B) models, in addition to LLaMA-3.1-8B. Table 3 reports performance for each model before and after fine-tuning with our method. Despite architectural and pretraining differences, Rewarding Doubt consistently reduces calibration error and improves AUROC across all models, without degrading downstream accuracy.
Table 3: Calibration and accuracy of Verbalize vs. Rewarding Doubt across different LLMs with 95% CIs in brackets.
Model Method ECE ($\downarrow$) AUROC ($\uparrow$) Accuracy ($\uparrow$)
LLaMA-3.1-8B Verbalize 0.2771 [0.2689,0.2862] 0.6766 [0.6667,0.6863] 0.6662 [0.6577,0.6745]
Trained probe 0.0152 [0.0118,0.0235] 0.8495 [0.8420,0.8567] 0.6231 [0.6143,0.6322]
Rew. Doubt 0.0256 [0.0209,0.0327] 0.8793 [0.8729,0.8860] 0.6497 [0.6407,0.6585]
Qwen-2.5-3B Verbalize 0.5330 [0.5252,0.5435] 0.5981 [0.5927,0.6035] 0.4185 [0.4085,0.4247]
Trained probe 0.0186 [0.0134,0.0268] 0.7975 [0.7880,0.8066] 0.2540 [0.2463,0.2624]
Rew. Doubt 0.1483 [0.1415,0.1546] 0.9065 [0.9012,0.9122] 0.4193 [0.4097,0.4283]
Qwen-2.5-7B Verbalize 0.3619 [0.3530,0.3705] 0.5818 [0.5762,0.5879] 0.5239 [0.5148,0.5331]
Trained probe 0.0989 [0.0920,0.1057] 0.8737 [0.8676,0.8797] 0.4793 [0.4696,0.4881]
Rew. Doubt 0.1298 [0.1237,0.1368] 0.8928 [0.8866,0.8988] 0.5283 [0.5193,0.5368]
Gemma-2-9B Verbalize 0.3206 [0.3122,0.3288] 0.5615 [0.5548,0.5682] 0.6690 [0.6603,0.6773]
Trained probe 0.0301 [0.0253,0.0373] 0.8694 [0.8629,0.8769] 0.6464 [0.6380,0.6551]
Rew. Doubt 0.0922 [0.0861,0.0994] 0.8649 [0.8570,0.8725] 0.6832 [0.6743,0.6918]
@@
To assess the generalization abilities of Rewarding Doubt, we evaluated the model trained on TriviaQA in out-of-domain settings using the CommonsenseQA (Talmor et al., 2019) and MedQA (Jin et al., 2020) datasets. Results are shown in Table˜4. On MedQA, Rewarding Doubt significantly outperforms Verbalize in both metrics, while on CommonsenseQA, it achieves a comparable ECE, however paired with a much higher AUROC. This discrepancy highlights a limitation of relying solely on ECE for evaluating calibration. ECE does not reflect how well a model discriminates between correct and incorrect predictions across different confidence levels. A model consistently assigning moderate confidence values could appear well-calibrated under ECE, yet fail to offer meaningful distinctions between uncertain and certain cases. AUROC, in contrast, directly measures this discriminative ability. Thus, the substantial improvements in AUROC underscore that Rewarding Doubt produces more useful and actionable confidence estimates. Compared to the Trained Probe, the best-performing baseline, Rewarding Doubt consistently outperforms, showing a stronger ability to generalize to new datasets.
Table 4: Comparison of generalization results on CommonsenseQA (CsQA) and MedQA, trained on the TriviaQA dataset with 95% CIs in brackets.
@@
CsQA Verbalize 0.2820 [0.2206,0.3422] 0.5425 [0.4740,0.6069] 0.6860 [0.6277,0.7444]
Trained Probe 0.4819 [0.4655,0.5130] 0.5374 [0.5021,0.5708] 0.7108 [0.6847,0.7355]
Rewarding doubt (ours) 0.2930 [0.2693,0.3179] 0.6385 [0.6065,0.6715] 0.7163 [0.6918,0.7410]
MedQA Verbalize 0.4480 [0.4200,0.4753] 0.5075 [0.4803,0.5338] 0.5067 [0.4784,0.5350]
Trained Probe 0.2099 [0.1881,0.2439] 0.5513 [0.5207,0.5844] 0.5051 [0.4792,0.5318]
Rewarding doubt (ours) 0.1145 [0.0893,0.1408] 0.6649 [0.6355,0.6954] 0.5161 [0.4886,0.5420]
We also explore generalization across experimental settings in Table˜5 by applying a model trained in a Single-Answer setting to a Multiple-Answer task. Although under-performing a model trained specifically for that task, it still outperforms the base model considerably, demonstrating transferability of the learned confidence estimation patterns. This suggests promising applications for improving confidence estimation in more complex or less structured scenarios, such as fact verification and calibration in free-text generation, even when specialized training data is unavailable.
Our current experiments focus on settings where answer quality can be evaluated via exact rule-based metrics, yielding a binary correctness signal, the Rewarding Doubt framework could be extended to work with correctness signals provided by an LLM-as-a-judge system, a reward model trained on human preferences or continuous NLG metrics.
Overall, our experiments show that Rewarding Doubt provides a robust and efficient way to enhance calibration, while generalizing across tasks, and maintaining stable task performance, making it an effective approach for accurate confidence calibration and expression in LLMs. Beyond improvements in calibration, our method also offers practical advantages. While fine-tuning requires an initial training investment, inference remains highly efficient, as only a small, constant number of tokens need to be generated to express confidence. In contrast, zero-shot methods like Chain-of-Thought and Self-Consistency have substantial computational overhead during inference by requiring lengthy reasoning chains or multiple generations. Rewarding Doubt introduces no such overhead, does not rely on an additional model, and directly provides actionable confidence estimates through verbalization directly by the LLM, making it highly suitable for real-world deployment.
Table 5: Comparison of the base and fine-tuned model on the Qampari dataset in different settings with 95% CIs in brackets.
@@
Base model Single fact 0.5875 [0.5597,0.6151] 0.5787 [0.5408,0.6125]
Single fact Single fact 0.1536 [0.1320,0.1813] 0.7240 [0.6889,0.7577]
Base model Multi fact 0.5319 [0.5172,0.5461] 0.6047 [0.5837,0.6267]
Single fact Multi fact 0.1777 [0.1679,0.1890] 0.6617 [0.6468,0.6779]
Multi fact Multi fact 0.1061 [0.0935,0.1206] 0.7268 [0.7065,0.7468]
@@
Due to computational constraints, we only tested Rewarding Doubt on models with sizes ranging from 3B to 9B parameters. While we expect similar effectiveness on larger models, empirical validation on such scales would be valuable.
@@ Acknowledgements
Reproducibility Statement
In order to ensure reproducibility, we describe implementation details of Rewarding Doubt as well as the used baselines in Section˜4 and Appendix˜D. Further, Appendix˜A provides the exact prompts used for different experiments. Lastly, we publish our code at https://github.com/pasta99/RewardingDoubt.
Appendix
@@ Appendix A Prompts
For all the question-answering settings, the model is directly prompted to answer a question without a preceding example or context. For our method the model was prompted to answer the question and additionally provide a verbalized confidence. For the other baselines that do not need a verbalized confidence but infer it indirectly, the model is prompted to only give the correct answer. The specifics for multiple-choice are slightly changed but hold mostly the same meaning. The exact prompts for each method can be seen in Table˜6 for open questions and Table˜7 for multiple-choice questions. The prompts for each Multi-Answer method can be seen in Table˜8. We decided not to give the model a role like "expert" in the system prompt but keep it neutral, as we observed that the role we give the model affects the verbalized confidence.
@@ Appendix B Proof
[…]
@@
is strictly negative for $\hat{p}\in(0,1)$. Hence, f(p^)f($\hat${p}) is concave and has its global maximum at $\hat{p}=p^{*}$. ∎
@@
[…]
@@
Appendix C Clipped Reward Function
The clipped reward function as described in Section˜3, is defined as follows:
[…]
where $\epsilon>0$ is a small positive constant of 0.001 introduced for numerical stability to avoid evaluating the logarithm at zero.
Appendix D Implementation Details of Baselines
For the Sequence Probability, we compute the average probability for each token in the response. In the Self-Consistency method, we let the model explore ten reasoning pathways, and the similarity of each resulting output is evaluated using the BERTScore metric Zhang et al. (2019). For the trained probe Azaria & Mitchell (2023), the original study introduced a custom dataset comprising short statements classified as either true or false. The model’s activations in response to these statements were extracted from specific layers, and a multilayer perceptron (MLP) was subsequently trained on these activations to predict the truthfulness of the statements. To ensure a fair comparison, we adapted this methodology to better align with our data by allowing the model to generate answers to training dataset questions and then extracting its activations from the 24th layer for both the statements and their corresponding answers. The labels for each sample were determined following the same evaluation procedure as described in our evaluation framework. For the architecture of the MLP, we employed the same design as Azaria & Mitchell (2023) and train it for four epochs with a learning rate of 1e-4 until convergence. The exact prompts used for each baseline are provided in Appendix˜A.
Appendix E Societal Impact
@@
Appendix F Use of Large Language Models
We employed ChatGPT to enhance the clarity of the manuscript by focusing on grammar corrections, shortening overly complex sentences, and providing alternative wording suggestions. All outputs were manually reviewed before inclusion, and no new technical material, code, results, or figures were generated by the tool.